OpenStudy (anonymous):

Double integral surface area:f(x,y)=(2/3)x^{3/2}+cos(x) 0

OpenStudy (anonymous):



OpenStudy (anonymous):

When the z function creates a boundary, does it define the region?

OpenStudy (turingtest):

Is this not a straight double integral then, it's a surface area?

OpenStudy (anonymous):

The region is given as a box however the f(x,y) splits the region in half

OpenStudy (anonymous):

surface area double integration

OpenStudy (turingtest):

sorry I'm too tired for this one now... next time g'night :)

OpenStudy (anonymous):



OpenStudy (sumeet):

is it that some part of volume is xy plane and some below it???

OpenStudy (anonymous):

or would it just be\[\int\limits_{0}^{1}\int\limits_{0}^{1}\frac{2}{3}x^{3/2}+\cos(x)\]

OpenStudy (sumeet):

yes, do it, you'll get a signed volume.

OpenStudy (anonymous):

that's for volume but for surface area

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{1}\sqrt{1+f_ x^2+f_ y^2}\]

OpenStudy (anonymous):

because for surface area you need a disk or region but it's giving the region already however i'd like to know what this looks like

OpenStudy (anonymous):

the question asks to use a computer tosolve it so the integrand is going to be horrible

OpenStudy (zarkon):

1.02185199655

OpenStudy (anonymous):

zarkon what would the surfacelook like?

OpenStudy (zarkon):

not pretty

OpenStudy (turingtest):

I'm glad I opted out.

OpenStudy (zarkon):

http://alturl.com/f5yz8

OpenStudy (sumeet):

true.. there's a way only for getting surface area for surface of revolutions. However what do you use for surface area of such equations.

OpenStudy (anonymous):

i was assuming it looked something like a plane

OpenStudy (turingtest):

No outkasts formula is more general than those for surface revolutions

OpenStudy (turingtest):

and you seem to be partly right :/

OpenStudy (anonymous):

partly right?

OpenStudy (turingtest):

I mean it looks almost a flat plane, a little curvier though.